At cryptobank a $100,000 loan for one year has a quoted interest rate of 10.50% payable annually but with a compensating balance of 20%. At Mo bank a $100,000 discount loan has a quoted interest rate of 7.90%payable annually. The loan at Mo bank is _____ per year cheaper than at cryptobank

1
Expert's answer
2022-04-17T17:09:33-0400

0.1050 "\cdot 0.2 =" 0.021

0.1050 - 0.021 = 0.084


0.084 - 0.079 = 0.005

by 5%